3/2y + 7/2=1/2y-1/2
Answer by gina(6)

3/2y + 7/2 = 1/2y - 1/2
3/2y - 1/2y + 7/2 = 1/2y - 1/2y - 1/2
y + 7/2 - 7/2 = -1/2 - 7/2
y = -8/2
y = -4